OpenStudy (ilovecake):

What is the LCD of 5/6 and -3/8 ?

OpenStudy (igreen):

Least Common Denominator. 5/6 Denominator = 6 -3/8 Denominator = 8 What's the smallest number 6 and 8 can go into?

OpenStudy (ilovecake):

I would say 18

OpenStudy (ilovecake):

24 I mean

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

6 * 4 = 24 8 * 3 = 24

OpenStudy (ilovecake):

ok thank you